A space traveler weighs 540.0 N on earth. Calculate what the travelers will weigh on another planet whose radius is four that of the earth and whose mass is five times that of earth ?
1
Expert's answer
2020-02-24T10:41:12-0500

Let's write the formula for the gravitational acceleration on Earth:


"g_E = \dfrac{GM_E}{R_E^2},"

here, "G" is the gravitational constant, "M_E" is the mass of the Earth, "R_E" is the radius of the Earth.

Then, we can write the gravitational acceleration on another planet in terms of "g_E" :


"g_{p} = \dfrac{GM_p}{R_p^2} = \dfrac{5GM_E}{(4R_E)^2} = \dfrac{5GM_E}{16R_E^2} = \dfrac{5}{16}g_E."

Finally, we can find the weight of the space traveler:


"W_{traveler, p} = mg_p = \dfrac{5}{16}mg_E = \dfrac{5}{16}W_{traveler,E} = \dfrac{5}{16} \cdot 540N = 169N."

Answer:

"W_{traveler, p} = 169N."